# On the Emergence of Position Bias in Transformers

### v3

@@ 3 Problem Setup
We use the shorthand $[n]:=\{1,\ldots,n\}$. For a matrix $M$, we denote its $i$-th row by $M_{i,:}$ and its $j$-th column by $M_{:,j}$. Throughout the analysis in the paper, we formalize the attention mask to be a directed graph $\mathcal{G}$. Formally, we represent a directed graph with $N$ nodes by $\mathcal{G}$ and let $E(\mathcal{G})$ be the set of directed edges of $\mathcal{G}$. A directed edge $(j,i)\in E(\mathcal{G})$ from node $j$ to $i$ in $\mathcal{G}$ means that in the attention mechanism, token $j$ serves as a direct context for token $i$ or token $i$ attends to token $j$. The set $\mathcal{N}_{i}$ of all neighbors of node $i$ is then $\{k:(k,i)\in E(\mathcal{G})\}$. We say a node $v$ is reachable from node $u$ in a directed graph $\mathcal{G}$ if there is a directed path $(u,n_{1}),(n_{1},n_{2}),...,(n_{k},v)$ from $u$ to $v$. In the attention mechanism, this means that token $u$ serves as a direct or indirect context for token $v$.
@@
Given the representation $X\in\mathbb{R}^{N\times d}$ of $N$ tokens, the raw attention score matrix is computed as $\small Z=XW_{Q}(XW_{K})^{\top}/\sqrt{d_{QK}}$, where $W_{Q},W_{K}\in\mathbb{R}^{d\times d^{\prime}}$ are the query and the key matrix, respectively, and $\sqrt{d_{QK}}$ is a temperature term to control the scale of raw attention scores. Without loss of generality, we assume $d_{QK}=1$ in our analysis. To enforce a masked attention, we create a sparse attention matrix $A\in\mathbb{R}^{N\times N}$ based on $Z$ whose sparsity pattern is specified by a directed graph $\mathcal{G}$: we normalize $Z_{ij}$ among all allowed token attention interactions $(k,i)\in E(\mathcal{G})$ such that if $(j,i)\in E(\mathcal{G})$,
@@ 4 Main Results
$X_{i,:}^{(t+1)}$ $=\sum_{j=1}^{N}\underbrace{(A^{(t)}\cdots A^{(0)})_{ij}}_{\mathclap{\mathbb{P}^{(t)}(z_{i}=j\mid X^{(0)})}}\,\cdot\,\underbrace{X^{(0)}_{j,:}W_{V}^{(0)}\cdots W_{V}^{(t)}}_{\mathclap{f^{(t)}(X^{(0)}_{z_{i},:})}},\vspace{-3ex}$ (4)
where $z_{i}$ is a categorical latent variable with a sample space $\{1,\ldots,N\}$ that selects the input $X_{j,:}$ to provide context for token $i$. In this formulation, $A^{(t)}$ represents the attention matrix at layer $t$, $\mathbb{P}^{(t)}(z_{i}=j\mid X^{(0)})$ denotes the cumulative probability of selecting input token $j$ as the context for token $i$ at depth $t$ , and $f^{(t)}(\cdot)$ is a learned transformation function.
@@ 4.1 Attention Masks: A Graph-Theoretic View
Moreover, there exist $10<C,\epsilon<$ 1 where $N\epsilon<1$ such that
@@
Moreover, there exist $10<C,\epsilon<$ 1 where $N\epsilon^{\left\lceil\frac{N-1}{w-1}\right\rceil}<1$ such that
@@
The above result suggests that a smaller window size $w$ helps mitigate the model’s bias toward early tokens in the sequence. However, such a moderating effect has its limit – the contextual information will still exponentially converge toward the first token over successive layers, though at a rate determined by the ratio between the sequence length $N$ and the window size $w$.
@@
Moreover, there exist $10<C,\epsilon<$ 1 where $N\epsilon<1$ such that
@@ 5 Experiments
Figure 2: Position bias arising solely from the architectural design of the attention mechanism, with no positional bias in the training data. $a$ vs. $b$ denotes the gap for the case $[\bm{a},b]-[\bm{b},a]$, where bar magnitude indicates gap size, positive indicates bias toward earlier position, and negative indicates bias toward later position. Deeper attention amplifies the bias toward earlier tokens, regardless of the PE used. Furthermore, decay mask introduce stronger distance-based decay effects that increase focus on recent tokens than RoPE. Residual connections affect positional bias in a nontrivial, depth- and PE-dependent way (see Section K.2 for more details).
@@ 5.1 The Effects of Depth and Relative PEs
To evaluate the position bias in the trained model, we construct test sequences of the form $[\bm{a},b]$. Here, the bolded term $\bm{a}$ explicitly marks the correct position, ensuring $y_{a}$ matches $y_{\text{query}}$, while position $b$ serves as a baseline. In these sequences, $x_{a}$ and $x_{b}$ are identical vectors, allowing us to control for the influence of semantic information on the model’s retrieval accuracy. We then measure the retrieval accuracy gap between pairs of sequences where the content at positions $a$ and $b$ is identical, but the correct position varies. This gap, defined as $[\bm{a},b]-[\bm{b},a]$, quantifies the model’s positional preference independent of semantic information. To perform this evaluation, we construct three pairs of test sets, each containing $10,000$ sequences: [first, middle] vs. [middle, first], [first, last] vs. [last, first], and [middle, last] vs. [last, middle]. Here “first” (position 1), “middle” (position $n/2$), and “last” (position $n$) denote fixed positions within a sequence.
Figure 2 shows the average results over five runs, where $a$ vs. $b$ denotes the gap $[\bm{a},b]-[\bm{b},a]$. The magnitude of each bar represents the size of the performance gap, and the sign of each bar reflects the direction of the bias: a positive sign indicates a bias toward earlier positions, while a negative sign indicates a bias toward later positions. We highlight several key observations. First, without residual connections, increasing model depth consistently amplifies the bias toward earlier parts of the sequence, regardless of the PE used (see Section K.1 for additional results). Also note that the performance gap between the middle and last positions is notably smaller than that involving the first position. This aligns with our theory, which suggests that as the model focuses more on the initial part of the sequence, information near the sequence’s end becomes less distinguishable, consistent with the patterns observed in Barbero et al. (2024a). Furthermore, both the decay mask and RoPE introduce distance-based decay effects that reduce the bias toward the beginning induced by the causal mask and increase the focus on recent tokens. However, the decay effect induced by the decay mask is substantially more pronounced than that by RoPE, as predicted by our theory.
@@ 5.2 Can Causal Mask Induce Usable Positional Information?
Figure 3 shows the average results using a $2$-layer network over five runs. Notably, in the left subplot, the causal mask without PE demonstrates a clear position bias toward the first position compared to the no mask without PE. This indicates that the causal mask indeed introduces a notion of position. However, when strong positional biases are present in the training data, both sin PE and RoPE allow the model to effectively capture these biases at both ends, regardless of the mask used. In contrast, a causal mask without PE only enables the model to learn a position bias at the beginning of the sequence. If the hypothesis by Kazemnejad et al. (2023) were correct, that the causal mask uses positional information by simulating PEs, then the model should be able to capture positional bias at any location. This discrepancy suggests that the causal mask does not inherently implement PE but instead introduces a bias toward earlier positions via iterative attention, capturing positional bias only when it aligns with this predisposition.
@@ A.1 Auxiliary results
We denote $P^{(t)}\vcentcolon=A^{(t)}\cdots A^{(0)}$. It suffices to show that there exists $10<C<$ 1 and $10<\epsilon<$ 1 such that
@@
Without loss of generality, we assume in the following proof that $N-1$ can be divided by $w-1$.
@@ B.1 Auxiliary results
Then by Lemma B.1 and Theorem 4.1, we get that there exists $10<C<$ 1 and $10<c<$ 1 such that for all $k\geq 0$
@@
Note that in the causal graph $\mathcal{G}$, there are ($t+i-j\choose i-j$ paths of length $t+1$ from token $j$ to token $i$.
@@
Given that when $\mathcal{G}$ is the causal graph, due to the connectivity the directed path of length $t$ from token $j$ to token $i$ must be non-decreasing, i.e. $j\leq l_{1}\leq l_{2}\leq\cdots\leq l_{t-1}\leq i$, and there would be in total (${t+i-j\choose i-j}$ such paths. For each such path $j\leq l_{1}\leq l_{2}\leq\cdots\leq l_{t-1}\leq i$, notice that by Lemma 4.6, we get that fix $T\geq 0$, there exists $C_{\min}$, $C_{\max}>0$ such that
@@ I.1 Results
In all experiments, we used the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of $10^{-3}$, a weight decay of $10^{-6}$, a batch size of $128$, and trained for $100,000$ iterations.

### v4

@@ 3 Problem Setup
We use the shorthand $n]:=\{1,\ldots,n\}$. For a matrix $M$, we denote its $i$-th row by $M_{i,:}$ and its $j$-th column by $M_{:,j}$. Throughout the analysis in the paper, we formalize the attention mask to be a directed graph $\mathcal{G}$. Formally, we represent a directed graph with NN nodes by $\mathcal{G}$ and let $E(\mathcal{G})$ be the set of directed edges of $\mathcal{G}$. A directed edge $(j,i)\in E(\mathcal{G})$ from node $j$ to $i$ in $\mathcal{G}$ means that in the attention mechanism, token $j$ serves as a direct context for token $i$ or token $i$ attends to token $j$. The set $\mathcal{N}_{i}$ of all neighbors of node $i$ is then $\{k:(k,i)\in E(\mathcal{G})\}$. We say a node $v$ is reachable from node $u$ in a directed graph $\mathcal{G}$ if there is a directed path $(u,n_{1}),(n_{1},n_{2}),...,(n_{k},v)$ from $u$ to $v$. In the attention mechanism, this means that token $u$ serves as a direct or indirect context for token $v$.
@@
Given the representation $X\in\mathbb{R}^{N\times d}$ of NN tokens, the raw attention score matrix is computed as $\small Z=XW_{Q}(XW_{K})^{\top}/\sqrt{d_{QK}}$, where $W_{Q},W_{K}\in\mathbb{R}^{d\times d^{\prime}}$ are the query and the key matrix, respectively, and $\sqrt{d_{QK}}$ is a temperature term to control the scale of raw attention scores. Without loss of generality, we assume $d_{QK}=1$ in our analysis. To enforce a masked attention, we create a sparse attention matrix $A\in\mathbb{R}^{N\times N}$ based on $Z$ whose sparsity pattern is specified by a directed graph $\mathcal{G}$: we normalize $Z_{ij}$ among all allowed token attention interactions $(k,i)\in E(\mathcal{G})$ such that if $(j,i)\in E(\mathcal{G})$,
@@ 4 Main Results
[…]
where $z_{i}$ is a categorical latent variable with a sample space $\{1,\ldots,N\}$ that selects the input $X_{j,:}$ to provide context for token ii. In this formulation, $A^{(t)}$ represents the attention matrix at layer $t$, $\mathbb{P}^{(t)}(z_{i}=j\mid X^{(0)})$ denotes the cumulative probability of selecting input token $j$ as the context for token $i$ at depth $t$ , and $f^{(t)}(\cdot)$ is a learned transformation function.
@@ 4.1 Attention Masks: A Graph-Theoretic View
Moreover, there exist $0<C,\epsilon<1$ where $N\epsilon<1$ such that
@@
Moreover, there exist $0<C,\epsilon<1$ where $N\epsilon^{\left\lceil\frac{N-1}{w-1}\right\rceil}<1$ such that
@@
The above result suggests that a smaller window size $w$ helps mitigate the model’s bias toward early tokens in the sequence. However, such a moderating effect has its limit – the contextual information will still exponentially converge toward the first token over successive layers, though at a rate determined by the ratio between the sequence length NN and the window size $w$.
@@
Moreover, there exist $0<C,\epsilon<1$ where $N\epsilon<1$ such that
@@ 5 Experiments
Figure 2: Position bias arising solely from the architectural design of the attention mechanism, with no positional bias in the training data. $a$ vs. $b$ denotes the gap for the case $\bm{a},b]-[\bm{b},a]$, where bar magnitude indicates gap size, positive indicates bias toward earlier position, and negative indicates bias toward later position. Deeper attention amplifies the bias toward earlier tokens, regardless of the PE used. Furthermore, decay mask introduce stronger distance-based decay effects that increase focus on recent tokens than RoPE. Residual connections affect positional bias in a nontrivial, depth- and PE-dependent way (see Section K.2 for more details).
@@ 5.1 The Effects of Depth and Relative PEs
To evaluate the position bias in the trained model, we construct test sequences of the form $\bm{a},b]$. Here, the bolded term $\bm{a}$ explicitly marks the correct position, ensuring $y_{a}$ matches $y_{\text{query}}$, while position $b$ serves as a baseline. In these sequences, $x_{a}$ and $x_{b}$ are identical vectors, allowing us to control for the influence of semantic information on the model’s retrieval accuracy. We then measure the retrieval accuracy gap between pairs of sequences where the content at positions $a$ and $b$ is identical, but the correct position varies. This gap, defined as $\bm{a},b]-[\bm{b},a]$, quantifies the model’s positional preference independent of semantic information. To perform this evaluation, we construct three pairs of test sets, each containing $10,000$ sequences: [first, middle] vs. [middle, first], [first, last] vs. [last, first], and [middle, last] vs. [last, middle]. Here “first” (position 1), “middle” (position $n/2$), and “last” (position $n$) denote fixed positions within a sequence.
Figure 2 shows the average results over five runs, where $a$ vs. $b$ denotes the gap $\bm{a},b]-[\bm{b},a]$. The magnitude of each bar represents the size of the performance gap, and the sign of each bar reflects the direction of the bias: a positive sign indicates a bias toward earlier positions, while a negative sign indicates a bias toward later positions. We highlight several key observations. First, without residual connections, increasing model depth consistently amplifies the bias toward earlier parts of the sequence, regardless of the PE used (see Section K.1 for additional results). Also note that the performance gap between the middle and last positions is notably smaller than that involving the first position. This aligns with our theory, which suggests that as the model focuses more on the initial part of the sequence, information near the sequence’s end becomes less distinguishable, consistent with the patterns observed in Barbero et al. (2024a). Furthermore, both the decay mask and RoPE introduce distance-based decay effects that reduce the bias toward the beginning induced by the causal mask and increase the focus on recent tokens. However, the decay effect induced by the decay mask is substantially more pronounced than that by RoPE, as predicted by our theory.
@@ 5.2 Can Causal Mask Induce Usable Positional Information?
Figure 3 shows the average results using a 22-layer network over five runs. Notably, in the left subplot, the causal mask without PE demonstrates a clear position bias toward the first position compared to the no mask without PE. This indicates that the causal mask indeed introduces a notion of position. However, when strong positional biases are present in the training data, both sin PE and RoPE allow the model to effectively capture these biases at both ends, regardless of the mask used. In contrast, a causal mask without PE only enables the model to learn a position bias at the beginning of the sequence. If the hypothesis by Kazemnejad et al. (2023) were correct, that the causal mask uses positional information by simulating PEs, then the model should be able to capture positional bias at any location. This discrepancy suggests that the causal mask does not inherently implement PE but instead introduces a bias toward earlier positions via iterative attention, capturing positional bias only when it aligns with this predisposition.
@@ A.1 Auxiliary results
We denote $P^{(t)}\vcentcolon=A^{(t)}\cdots A^{(0)}$. It suffices to show that there exists $0<C<1$ and $0<\epsilon<1$ such that
@@
Without loss of generality, we assume in the following proof that N1N-1 can be divided by w1w-1.
@@ B.1 Auxiliary results
Then by Lemma B.1 and Theorem 4.1, we get that there exists $0<C<1$ and $0<c<1$ such that for all $k\geq 0$
@@
Note that in the causal graph $\mathcal{G}$, there are $t+i-j\choose i-j$ paths of length $t+1$ from token $j$ to token ii.
@@
Given that when $\mathcal{G}$ is the causal graph, due to the connectivity the directed path of length $t$ from token $j$ to token $i$ must be non-decreasing, i.e. $j\leq l_{1}\leq l_{2}\leq\cdots\leq l_{t-1}\leq i$, and there would be in total $t+i-j\choose i-j$} such paths. For each such path $j\leq l_{1}\leq l_{2}\leq\cdots\leq l_{t-1}\leq i$, notice that by Lemma 4.6, we get that fix $T\geq 0$, there exists $C_{\min}$, $C_{\max}>0$ such that
@@ I.1 Results
In all experiments, we used the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of $10^{-3}$, a weight decay of $10^{-6}$, a batch size of 128128, and trained for $100,000$ iterations.